What is the average kinetic energy of a molecule at (a) 0°C and (b) 500°c? [5.66 x 10.21 12.64 x 1020 1)
Answer #1

ANSWER:

Boltzman constant, KB = 1.38 x 10-23 J/K

(a) 0 oC

T = 0 oC = 273.15 K

At this temperature only translation degree of freedom is taken into consideration. So,

Av. kinetic energy of a molecule = (3/2)KBT

= (3/2) x 1.38 x 10-23 J/K x 273.15 K

= 5.66 x 10-21 J

Hence, the average kinetic energy of a molecule at 0 oC is 5.66 x 10-21 J.

(b) 500 oC

T = 500 oC = 773.15 K

At this temperature rotational and translation degree of freedom both taken into the consideration. So,

Av. kinetic energy of a molecule = (5/2)KBT

= (5/2) x 1.38 x 10-23 J/K x 773.15 K

= 2.66 x 10-20 J

Hence, the average kinetic energy of a molecule at 500 oC is 2.66 x 10-20 J.